A bag contains 6 cherry, 3 orange and 2 lemon candies. You reach in and take 3 pieces of candy at random. What is the probability that you have at least 2 orange candies?
Answer #1
number of ways to select 3 from 11 ___ 11C3 = 165 ways to select 2 orange out of 3 ___ 3C2 = 3 ways to select 2 orange out of 11 ___ 3 * 8 = 24 ways to select 3 orange out of 11 ___ 1 probability of at least 2 orange ___ (24 + 1) / 165 = 5 / 33
answered by: Doret